Jane planted a rectangular garden that is 12 feet wide by 17 feet long. What is the perimeter of jane's garden? Answer in units of ft.
Answer by ikleyn(52780)

.
12 + 17 + 12 + 17 = 2*(12 + 17) = 2*29 = 58 ft.